OpenStudy (anonymous):

a particular sound wave can be graphed using the function y = -3 sin x. find the period of the function.

OpenStudy (anonymous):

All that's happened is the sine function graph has been stretched vertically by 3 and reflected about the x-axis. The period is identical to that of the parent function \(\sin x\). Do you know the period of \(\sin x\)?

OpenStudy (anonymous):

Amplitude is defined by y=a*sin(b*x+c)+d where a=amplitude period is represented usually by a T and there T= 2pi/b the Phase shift is c/b http://www.teacherschoice.com.au/maths_library/functions/about_trigonometric_functions.htm Think you can now figure out what the amplitude is or do you still need some help?

OpenStudy (anonymous):

@FutureMathProfessor I don't need to know the amplitude, I need to know the period. Unless that still applies. @oldrin.bataku no i don't, there's not even a graph

OpenStudy (anonymous):

@meegan are you SURE you don't know the period of \(\sin x\)?

OpenStudy (anonymous):

It takes \(2\pi\) to go around the circle once and start over... the period of \(\sin x\) is thus \(2\pi\). The function you have here, \(y=-3\sin x\), is merely \(\sin x\) stretched vertically (hence the multiplication by \(3\)) and reflected about the \(x\)-axis (hence the \(-\) sign). Neither of these affect its horizontal behavior, so the function will have the same period as \(\sin x\).

OpenStudy (anonymous):

so the period is 2pi?

OpenStudy (anonymous):

or is it 3

OpenStudy (anonymous):

The period of \(\sin x\) is \(2\pi\). You can shift, scale it vertically, reflect it about the axes -- the period won't change. Hence the period of \(-3\sin x\) is also \(2\pi\).

OpenStudy (anonymous):

The period *can* change if you do something like \(\sin kx\), though!

OpenStudy (anonymous):

okay great. thanks!